OpenStudy (anonymous):

help plz solve:- (log X)^2 + log X^2 = (log2)^2 -1

OpenStudy (anonymous):

\[(\log X)^{2} + \log X ^{2} =(\log 2)^{2} -1\]

OpenStudy (anonymous):

what kind of problem is this

OpenStudy (anonymous):

solve find x = ?

OpenStudy (anonymous):

linear equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not sure can u find a calculator on google to do it?

OpenStudy (anonymous):

it so easy

OpenStudy (anonymous):

u got it already?

OpenStudy (anonymous):

I can't

OpenStudy (anonymous):

you put its sooo easy

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

can you solve IT ?

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

Using properties of logs rewrite it like this (log X)^2 + 2log X = (log2)^2 -1 Take the 1 to the other side (log X)^2 + 2log X +1 = (log2)^2 Notice the quadratic on the left (log(X)+1)^2=(log2)^2 Can you go from there?

OpenStudy (anonymous):

ok